Circular No. 7728

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001eq IN PGC 70417 W. D. Li, University of California at Berkeley, reports the discovery

by LOTOSS (cf. IAUC 7514) of an apparent supernova in unfiltered images taken with the Katzman Automatic Imaging Telescope. SN 2001eq is located at α = 23^{h}04^{m}56^{s}.78, δ = +19^{o}33′04′′.8 (equinox 2000.0), which is 3′′.0 east and 2′′.6 south of the nucleus of PGC 70417. Magnitudes for SN 2001eq: Aug. 23.3 UT, [19.0; 31.3, [18.5; Sept. 12.3, 18.7; 17.3, 18.0; 23.3, 17.6; Oct. 4.3, 18.5.

SUPERNOVA 2001er IN UGC 5301 T. Puckett, Mountain Town, GA; and D. George, Ottawa, ON, report

the discovery of an apparent supernova (mag 17.2) on an unfiltered CCD frame (limiting mag 18.8) taken with the Puckett Observatory 0.35-m automated supernova patrol telescope on Oct. 4.42 UT. SN 2001er is located at α = 9^{h}53^{m}20^{s}.66, δ = +42^{o}50′44′′.8 (equinox 2000.0), which is 5′′.0 west and 10′′.1 north of the center of UGC 5301. The new object was confirmed on CCD frames taken on Oct. 5.49 with a 0.40-m reflector by J. Newton, Osoyoos, BC. SN 2001er is not present on images taken by Puckett on 2000 Jan. 2 and 9, Mar. 26, and Apr. 25 (limiting mag ∼ 20.5), neither is it present on Palomar Sky Survey images taken on 1995 Mar. 28, 1991 Dec. 13 (limiting mag ∼ 21.0), and 1953 Feb. 8 (limiting mag ∼ 19.7).

V445 PUPPIS M. P. Rupen, A. J. Mioduszewski, and V. Dhawan, National Radio

Astronomy Observatory, write: “Observations with the Very Large Array show that V445 Pup is currently undergoing a renewed, strong radio flare. Data taken on Oct. 4 give flux densities of 11.72 ± 0.23, 14.20 ± 0.10, and 12.17 ± 0.08 mJy at 1.425, 4.860, and 8.460 GHz, respectively. The 4.86and 8.46-GHz flux densities are significantly higher than those measured on Sept. 25 (8.3 ± 0.1 and 7.1 ± 0.1 mJy, respectively). The 1.425-GHz measurement is much lower than a simple extrapolation of the high-frequency data would suggest, implying substantial absorption. The opacity inferred from earlier radio spectra was much lower, so we are probably seeing synchrotron self-absorption in a small (on the order of milliarcseconds) radio source that becomes optically thin as it expands. Further observations at other wavelengths are strongly encouraged.”

2001 October 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green